Calculate the energy stored in elastic material with K=1500NM and applied force at 750N

1
Expert's answer
2022-03-17T13:55:31-0400

Answer


Stored energy

"U=\frac{kx^2}{2}=\frac{F^2}{2K}\\=\frac{750*750}{3000}\\=187.5J"